The Magellanic Clouds Newsletter

An electronic source for Magellanic Clouds informations

Edited by: You-Hua Chu and Dominik J. Bomans

mcnews@astro.uiuc.edu

Number: 0 September 15, 1995

Dear colleague:

Greetings!

A group of us in the Midwest have been working on the Magellanic Clouds (MCs). We would like to start an electronic publication of the Magellanic Clouds Newsletter to promote information exchange and interdisciplinary collaboration. We plan to publish the newsletter bi-monthly or monthly (depending on the response from readers) in Latex format, and deliver the first issue on December 1, 1995. A WWW page for the newsletter is being constructed and will be introduced soon.

The Magellanic Clouds Newsletter will include the following:

• abstracts of accepted papers

• abstracts of submitted papers

• news

– reports of large surveys of the Magellanic Clouds

– information/data requests

– meeting announcements

– readers’ forum

In the ”news” section, we hope to stimulate interactions among readers by reporting the progress of surveys, posting requests for information or data, and announcing meetings. If there is sufficient interest, a controversial topic may be selected for the readers’ forum. (Contributions to the ”news” section can be submitted in either ascii or Latex format.)

Your name has been included in this mailing list because you have attended one of the meetings on the MCs or have worked on subjects related to the MCs. If you wish to remove your name from the list or add more names to the list, please send us an e-mail.

If you are interested in the MC Newsletter, please send us abstracts of your papers. A Latex template (adopted from the Hot Star Newsletter) is attached below. Suggestions and comments are welcome, too.

If you are hard-core fans of the MCs, please note a ”possible” special session on the MCs in the American Astronomical Society (AAS) meeting in Madison, Wisconsin on June 9-13, 1996. We are going to propose to the AAS for a special session dedicated to the MCs studies. We plan to request one and a half days for this session; 1/2 day for stars, 1/2 day for the interstellar medium, and 1/2 day for interactions between stars and the interstellar medium. Each 1/2 day will start with an overview talk, followed by reports of large surveys and individual contributions. If you have suggestions or would like to contribute, please answer the questionnaire attached at the end of this message.

Please send your correspondence via e-mails to

mcnews@astro.uiuc.edu

We look forward to your feedback. Thanks.

Best wishes,

You-Hua Chu and Dominik Bomans

Astronomy Department, University of Illinois

% WHEN SUBMITTING AN ABSTRACT, PLEASE USE THE FOLLOWING TEMPLATE
% AND SUBMIT IT TO     mcnews@astro.uiuc.edu
\begin{center}{\Large\bf  Title }\end{center}
\centerline{\bf         A. Author$^1$ and B. Author$^2$}
{\footnotesize   $^1$  Institute One and Address
         \\       $^2$  Institute Two and Address
}
\vspace*{4mm}\\
Text of abstract
{\bf   Accepted by
%          or
      Submitted to      JOURNAL
}
{\it For preprints, contact\,}              your electronic address
% (please also indicate if the manuscript is available by anonymous ftp)

*******************************************************************************
                           MAGELLANIC CLOUDS
     Possible Special Section at AAS meeting in Madison, Wisconsin
                    1.5 days during June 9-13, 1996
*******************************************************************************
Are you interested in attending:   ___yes     ___no      ___probably
Would you plan to to present contribution:    ___yes     ___no
   if yes, what will be the topic:
   _____________________________________________________________________
Suggestions:
  _____________________________________________________________________
  _____________________________________________________________________
Name: _________________________________________________________________
Institution: __________________________________________________________
E-mail address: _______________________________________________________
Please return this questionnaire to mcnews@astro.uiuc.edu by October 31.
We need your input to prepare the proposal the AAS.  Thanks.